OpenStudy (anonymous):

I'm so confused. Any help would be appreciated. sin(2cos^-1(7/25))

OpenStudy (anonymous):

\[\sin (2\cos^{-1}( \frac{ 7 }{ 25 }))\] Is this the equation?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

do you have a scientific calculator?

OpenStudy (anonymous):

I do

OpenStudy (anonymous):

find the value of the angle then multiply by two... then sin the value

OpenStudy (anonymous):

then just type the whole thing in cos-1 is shift+cos

OpenStudy (anonymous):

don't forget to use parenthesis in the calculator make sure it is in rad or deg... rad if you are looking for the values in radians and deg for degrees

OpenStudy (anonymous):

OK thanks all

OpenStudy (anonymous):

btw just for the image since i already uploaded http://i62.tinypic.com/fuvqbq.png

OpenStudy (anonymous):

Thanks Auxuris. Do you get .5376?

OpenStudy (anonymous):

yuuup

OpenStudy (anonymous):

ok thanks